Monday Operations Meeting : 23/3/2015¶
Agenda:¶
- Action Items from previous meeting
Show Dean the latest plan for the Wednesday tuning procedure and get comments (Steve)- Define BLOC cover (Henry/Ken/Steve)
Organise a SOC phone (Steve)- Check the progress of the action items from the Mock Data Runs with Pierrick and Rhys (Steve)
Contact Elaine Lancaster regarding MICE/ISIS meeting on Friday (Ken)
- Review of weekend run
- VERY PRELIMINARY TOF calibration:
- Trigger from TOF0 Tof0_resol_trtof0.png, Tof1_resol_trtof0.png ToF02ToF1.png
- Trigger from TOF1 Tof0_resol_trtof1.png, Tof1_resol_trtof1.png
- Final calibration to be made by Durga
- 2015-03-23 Physics shifter report
- VERY PRELIMINARY TOF calibration:
Decisions made ¶
- We will keep the run condition spreadsheet going in parallel with the CDB until we are sure that the
CDB contains all the information we need. - Be more methodical when going the checklists
- Actions from previous runs
- Fence from Q9 down to main door should be in place
- Cranes should be in right place
- Method statement for operating beamline not followed
- Overview of readiness list from MDR1 and MDR2 : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Operations_Readiness_To_Do_List
- Plans for March 28/29 running
- Data taking plans : http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20150328
- Training plans
- Shifters
Decisions made¶
- Check pion run from last weekend is OK (Victoria)
- Saturday data taking will be debug session to deal with the little problems that have been observed so far
- Sunday :
**Take data without trigger system for comparison
**Do CKOV HV scan
- At shift handover Henry will run extra target training
- Friday training
- 9:30 - 10:30 : BLOC Training
- 1 pm : Control room training
- 3 pm : Hall search. No PPS tutorial, but refresher for shifters.
- Longer term plans
- 28/29Mar15: CKOV, TOF0, TOF1, no DS
*Establish smooth operation from target to TOF1;
*Carry out an initial calibration of TOF0 and TOF1;
*Take data to allow a comparison of the performance of the old and new trigger systems;
*Take pion-beam data to allow a first test of the reconstruction algorithms required for MICE Muon Beam commissioning;
*Complete BLOC training on target tuning.
- 11/12Apr15, 18/19Apr15 and 25/26Apr15: DS, CKOV,TOF0,TOF1, EMR, KL, TOF2 (& tracker?)
*Check the TOF calibration. It will be important to take some data to be sure that the calibration is stable;
*First pass of MICE Muon Beam (MMB) commissioning:
*Start with the currents used in Step I to check the degree of reproducibility of particle rates;
*Based on the revised simulation of the beam line (Jaroslaw Pasternak), start the search for the optimum magnet settings;
*Commissioning and calibration of TOF2, KL and EMR will begin as soon as the south-side PRY is complete and the magnets and downstream
detectors are in their final Step IV positions.
- 28/29Mar15: CKOV, TOF0, TOF1, no DS
- AOB
- Next meeting : 11 am Friday 27th March
Phone access : http://www.mice.iit.edu/phonebridge.html
Actions arising from meeting¶
- Check with Andy on the safety implications of the change of hall configuration during search (Steve)
- minor changes to checklists (Steve)
- Save current run summary spreadsheet from Sunday run to wiki (?)
- Enhance Elog DAQ error messages so that physics shifter has a better idea of what went wrong and whether it is serious. Include error message in Elog entry (All)
- Think about adding an "Analysable" flag to the run conditions spreadsheet for the near term (Chris)
- Chase the OnMon changes in the MRD action items (Steve)
- General comment : Please be methodical when going through checklists and ensure that all steps are carried out. (All)
Updated by Boyd, Steven over 8 years ago ยท 16 revisions